The perimeter of a triangle is 124.
The first side is 6cm less than 3 times the second. The 3rd side is 10cm more than the second. what are the lengths of each side?

Answer by Sarah900(12)

To solve this problem we need to write the side lengths in terms of a single variable. Since all the side lengths are compared to the second side, we will call the second side's length "x". So...


Second length=x
first length= 3x-6
third length= x+10


Since the perimeter of a triangle is equal to the length of all three sides we can add the three..
124= x+(x+10)+(3x-6)
Combine like terms and solve for x
124=5x+4
120=5x
24=x


Now we know the length of the second side which we originally deemed x. To find the others you replace x with 24 in the other two.
3(24)-6=66
24+10=34